# Evaluating Memory in LLM Agents via Incremental Multi-Turn Interactions

### v1

@@ C.3 Detailed Results on Ablation Study
In Table 12 and Table 14, we provide the computational latency on different agents. We choose two chunk sizes {512, 4096} and two datasets {RULER-QA2, LME(S*)}. The experimental results demonstrate that, for most agents, selecting a smaller chunk size leads to significantly higher computational latency. For example, in the case of Cognee, the computational latency at a chunk size of 512 is nearly 8 to 10 times greater than that at a chunk size of 4096.
| NIAH-MQ | Bench-QA | LME(S*) | Event-QA | FactCon-SH | FactCon-MH | |||||||
| 512 | 4096 | 512 | 4096 | 512 | 4096 | 512 | 4096 | 512 | 4096 | 512 | 4096 | |
| Simple RAG Agents | ||||||||||||
| BM25 | 100 | 95.5 | 32.9 | 45.6 | 45.3 | 48.3 | 69.4 | 74.6 | 56.0 | 44.0 | 3.0 | 5.0 |
| Embedding RAG Agents | ||||||||||||
| Contriever | 2.5 | 8.8 | 28.5 | 38.1 | 15.7 | 19.0 | 62.8 | 66.8 | 18.0 | 25.0 | 7.0 | 5.0 |
| Text-Embed-3-Small | 7.2 | 12.3 | 42.4 | 44.4 | 48.3 | 39.0 | 60.8 | 63.0 | 28.0 | 21.0 | 3.0 | 4.0 |
| Text-Embed-3-Large | 19.5 | 13.5 | 42.3 | 50.1 | 52.3 | 39.3 | 62.0 | 70.0 | 28.0 | 22.0 | 4.0 | 3.0 |
| NV-Embed-v2 | 73.5 | 31.8 | 40.7 | 51.4 | 55.0 | 43.0 | 72.7 | 72.8 | 55.0 | 42.0 | 6.0 | 6.0 |
| Structure-Augmented RAG Agents | ||||||||||||
| RAPTOR | 15.8 | 4.5 | 21.3 | 31.3 | 34.3 | 31.7 | 43.6 | 45.8 | 14.0 | 19.0 | 1.0 | 2.0 |
| GraphRAG | 38.3 | 8.0 | 35.2 | 35.8 | 35.0 | 36.7 | 33.8 | 34.4 | 14.0 | 10.0 | 2.0 | 3.0 |
| HippoRAG-v2 | 67.5 | 23.3 | 34.5 | 45.7 | 50.7 | 37.3 | 67.5 | 67.6 | 54.0 | 29.0 | 5.0 | 3.0 |
| Agentic Memory Agents | ||||||||||||
| Self-RAG | 8.0 | 7.0 | 27.7 | 28.5 | 25.7 | 23.0 | 30.2 | 31.8 | 19.0 | 14.0 | 3.0 | 2.0 |
| MemGPT | 8.8 | 3.5 | 23.5 | 20.8 | 32.0 | 28.0 | 26.2 | 25.8 | 28.0 | 13.0 | 3.0 | 3.0 |
| RULER-QA-1 | RULER-QA-2 | Bench-Sum | ||||||||||
@@
| RULER | NIAH | Bench-QA | EventQA | TTL (MCC) | |||||||||||
| R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | |
| BM25 | 49.5 | 59.5 | 61.0 | 34.5 | 57.0 | 95.5 | 26.7 | 38.3 | 45.6 | 66.6 | 71.2 | 74.6 | 67.8 | 75.4 | 74.6 |
| Contriever | 26.0 | 38.0 | 41.0 | 0.5 | 5.3 | 8.8 | 23.8 | 37.1 | 38.1 | 54.4 | 66.8 | 56.0 | 63.0 | 70.0 | 70.6 |
| Text-Embed-3-Large | 32.5 | 33.5 | 36.5 | 5.0 | 9.3 | 13.5 | 34.8 | 41.9 | 50.1 | 51.8 | 62.4 | 70.0 | 59.4 | 69.4 | 72.4 |
| NV-Embed-v2 | 37.0 | 43.5 | 48.0 | 17.8 | 26.8 | 31.8 | 42.8 | 48.1 | 51.4 | 59.4 | 68.4 | 72.8 | 63.8 | 69.4 | 68.8 |
| RAPTOR | 21.0 | 19.5 | 23.5 | 4.3 | 4.5 | 4.3 | 30.9 | 30.4 | 31.3 | 45.8 | 41.8 | 40.4 | 56.3 | 59.4 | 57.4 |
| HippoRAG-v2 | 38.0 | 42.5 | 43.5 | 16.5 | 23.3 | 23.3 | 35.9 | 45.3 | 45.7 | 58.8 | 67.6 | 67.4 | 58.8 | 61.4 | 61.4 |
| Self-RAG | 29.5 | 32.5 | 38.5 | 4.0 | 6.3 | 7.0 | 21.0 | 23.9 | 28.5 | 28.2 | 30.6 | 31.8 | 9.0 | 11.6 | 11.6 |
| RULER | NIAH | EventQA | FactCon-SH | FactCon-MH | |||||||||||
| 51K | 104K | 304K | 55K | 117K | 448K | 51K | 108K | 534K | 32K | 64K | 262K | 32K | 64K | 262K | |
| GPT-4o | 81.5 | 76.0 | 61.5 | 100 | 100 | 25.0 | 96.8 | 94.0 | 77.2 | 88.0 | 85.0 | 60.0 | 10.0 | 13.0 | 5.0 |
| GPT-4o-mini | 71.0 | 68.5 | 53.5 | 99.5 | 99.0 | 22.8 | 90.2 | 85.8 | 59.0 | 63.0 | 58.0 | 45.0 | 10.0 | 5.0 | 5.0 |
| GPT-4.1-mini | 82.5 | 80.5 | 74.5 | 99.5 | 99.0 | 94.8 | 97.0 | 93.8 | 82.6 | 82.0 | 72.0 | 36.0 | 7.0 | 9.0 | 5.0 |
| Gemini-2.0-Flash | 80.5 | 74.0 | 73.0 | 87.8 | 93.3 | 83.8 | 93.4 | 88.6 | 67.2 | 49.0 | 62.0 | 30.0 | 7.0 | 9.0 | 3.0 |
| Claude-3.7-Sonnet | 78.5 | 70.5 | 65.0 | 99.5 | 100 | 38.0 | 96.6 | 95.2 | 74.6 | 46.0 | 45.0 | 43.0 | 2.0 | 2.0 | 0.0 |
| Mem0 | 28.5 | 27.0 | 28.0 | 5.5 | 5.0 | 4.8 | 60.8 | 47.0 | 40.2 | 22.0 | 8.0 | 18.0 | 3.0 | 2.0 | 2.0 |
| Cognee | 37.0 | 40.0 | 33.5 | 17.5 | 14.3 | 4.0 | 53.4 | 39.0 | 26.8 | 39.0 | 31.0 | 28.0 | 4.0 | 5.0 | 3.0 |
| RULER-QA2 | LME (S*) | |
@@
Appendix D Experimental Settings
In this section, we present the experimental settings.
D.1 Max Output Tokens
We provide the token number limitation for each task in Table 14.
D.2 Settings of the RAG Agents
For the embedding model selection in Structure-Augmented RAG Agents and Agentic Memory Agents, most approaches utilize OpenAI’s embedding models, such as Text-Embed-3-Small. While for the HippoRAG-v2 method, we follow the same experimental setting as in Gutiérrez et al. [12], employing the NV-Embed-v2 model.
D.3 Settings of the Chunk Size
We use smaller chunk size (512) for synthetic context used in AR and CR. For some tasks based on continuous text, such as Bench and EventQA, we used a larger chunk size (4096). For tasks such as MCC, Recom and LME(S), considering the characteristics of these tasks and the computational cost, we also chose a larger chunk size (4096). For the two memory construction methods that are more time-consuming, Mem0 and Cognee, we uniformly used a chunk size of 4096 across all datasets.

### v2

@@ D.4 Results on Different Context Length Ablation Study
| SH-Doc QA | MH-Doc QA | Bench-Sum | ||||||||||
@@
| SH-Doc QA | MH-Doc QA | EventQA | TTL (MCC) | |||||||||
| R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | |
| BM25 | 50.0 | 60.0 | 66.0 | 49.0 | 54.0 | 56.0 | 66.6 | 71.2 | 74.6 | 67.8 | 74.6 | 75.4 |
| Contriever | 17.0 | 20.0 | 22.0 | 22.0 | 27.0 | 31.0 | 54.4 | 66.8 | 56.0 | 63.0 | 70.0 | 70.6 |
| Text-Embed-3-Large | 36.0 | 47.0 | 54.0 | 37.0 | 41.0 | 44.0 | 51.8 | 62.4 | 70.0 | 59.4 | 69.4 | 72.4 |
| RAPTOR | 22.0 | 27.0 | 29.0 | 30.0 | 36.0 | 38.0 | 45.8 | 41.8 | 40.4 | 56.3 | 57.4 | 59.4 |
| HippoRAG-v2 | 60.0 | 69.0 | 76.0 | 53.0 | 60.0 | 66.0 | 58.8 | 67.6 | 67.4 | 58.8 | 61.4 | 61.4 |
| Self-RAG | 27.0 | 33.0 | 35.0 | 34.0 | 39.0 | 42.0 | 28.2 | 30.6 | 31.8 | 9.0 | 11.6 | 11.6 |
| SH-Doc QA | MH-Doc QA | EventQA | FactCon-SH | FactCon-MH | |||||||||||
| 51K | 104K | 197K | 51K | 104K | 421K | 51K | 108K | 534K | 32K | 64K | 262K | 32K | 64K | 262K | |
| GPT-4o | 91.0 | 84.0 | 72.0 | 72.0 | 68.0 | 51.0 | 96.8 | 94.0 | 77.2 | 88.0 | 85.0 | 60.0 | 10.0 | 13.0 | 5.0 |
| GPT-4o-mini | 84.0 | 83.0 | 64.0 | 58.0 | 54.0 | 43.0 | 90.2 | 85.8 | 59.0 | 63.0 | 58.0 | 45.0 | 10.0 | 5.0 | 5.0 |
| GPT-4.1-mini | 93.0 | 86.0 | 83.0 | 72.0 | 75.0 | 66.0 | 97.0 | 93.8 | 82.6 | 82.0 | 72.0 | 36.0 | 7.0 | 9.0 | 5.0 |
| Gemini-2.0-Flash | 92.0 | 87.0 | 87.0 | 69.0 | 61.0 | 59.0 | 93.4 | 88.6 | 67.2 | 49.0 | 62.0 | 30.0 | 7.0 | 9.0 | 3.0 |
| Claude-3.7-Sonnet | 90.0 | 82.0 | 77.0 | 67.0 | 59.0 | 53.0 | 96.6 | 95.2 | 74.6 | 46.0 | 45.0 | 43.0 | 2.0 | 2.0 | 2.0 |
| Mem0 | 31.0 | 25.0 | 25.0 | 36.0 | 29.0 | 32.0 | 60.8 | 47.0 | 37.5 | 22.0 | 8.0 | 18.0 | 3.0 | 2.0 | 2.0 |
| Cognee | 38.0 | 42.0 | 31.0 | 36.0 | 38.0 | 26.0 | 53.4 | 39.0 | 26.8 | 39.0 | 31.0 | 28.0 | 4.0 | 5.0 | 3.0 |
D.5 Results on Computational Latency Analysis
To illustrate the latency of various memory agents in terms of (1) Memory Construction (M.C.); (2) Query Execution (Q.E.), we report the latency of various memory agents on MH-QA and LME (S*). This part of experiments is done on a server with Four NVDIA L40 GPU and AMD EPYC 7713 64-Core CPU. We use the NV-Embed-v2 (7B) as the embedding model in HippoRAG-v2. We show the results in Table 11 and 11. From the table, we find that using a smaller chunk size requires significantly more time for memory construction, especially for methods such as HippoRAG-v2, Mem0, Cognee, and MemGPT. Meanwhile, methods such as Mem0, Cognee and MIRIX need extremely high resources when constructing the memory.
| MH-QA | LME (S*) | |
@@ D.5 Results on Computational Latency Analysis
| MIRIX | 29000* | - | 20171 | 14.1 | 12600* | - | 3258 | 8.7 |
| MIRIX (GPT-4.1-mini) | 28800* | - | 21361 | 16.9 | 9000* | - | 2512 | 9.2 |
D.6 GPU Memory Usage Comparison
In main experiments, we mostly use the LLM API as the backbone models which do not need local GPUs. In our experiments, the HippoRAG-v2 (NV-Embed-v2) and Qwen3-Embedding-4B require running the embedding model on GPU. We report their peak GPU memory usage in Table 12, where all experiments are conducted on a single A100 80GB GPU.
| Agents / Chunk Size | 512 | 4096 |
| HippoRAG-v2 (NV-Embed-v2) | 27674 | 60205 |
| Qwen3-Embedding-4B | 16671 | 41262 |
Appendix E Experimental Settings
In this section, we present the experimental settings in evaluation.
E.1 Max Output Tokens
We provide the token number limitation for each task in Table 14.
E.2 Settings of the RAG Agents
For the embedding model selection in Structure-Augmented RAG Agents and Agentic Memory Agents, most approaches utilize OpenAI’s embedding models, such as Text-Embed-3-Small. While for the HippoRAG-v2 method, we follow the same experimental setting as in Gutiérrez et al. (2025), employing the NV-Embed-v2 model.
E.3 Settings of the Chunk Size
We use smaller chunk size (512) for synthetic context used in AR and SF. For some tasks based on continuous text, such as Bench and EventQA, we used a larger chunk size (4096). For tasks such as MCC and Recom, considering the characteristics of these tasks and the computational cost, we also chose a larger chunk size (4096). For the memory construction methods that are more time-consuming and requiring more API cost, Mem0, Zep, Cognee and MIRIX, we uniformly used a chunk size of 4096 across all datasets. The detailed settings are presented in Table 14.